Find taylor series
F(x) = (6-x)^-7 About x=4
1
Expert's answer
2020-10-08T16:16:01-0400

Given expression is "f(x) = \frac{1}{(6-x)^7}"

First derivative, "f'(x) = \frac{7}{(x-6)^8}"

Second derivative "f''(x) = -\frac{56}{(x-6)^9}"

Third derivative "f'''(x) = \frac{504}{(x-6)^{10}}"

Fourth derivative "f''''(x) = -\frac{5040}{(x-6)^{11}}"

and so on .........


Value of f(x) at x=4, "f(4) = \frac{1}{128}"

Value of first derivative at x=4, "f'(4) = \frac{7}{256}"

Value of second derivative at x=4, "f''(4) = \frac{7}{64}"

Value of third derivative at x=4, "f'''(4) = \frac{63}{128}"

Value of fourth derivative at x=4, "f''''(4) = \frac{315}{128}"


Taylor series of f(x)


"f(x) = f(4) + \frac{f'(4)}{1!}(x-4) + \frac{f''(4)}{2!}(x-4)^2 + \frac{f'''(4)}{3!}(x-4)^3 + \frac{f''''(4)}{4!}(x-4)^4+ ....."

"f(x) = \frac{1}{128} + \frac{7}{256}(x-4) + \frac{7}{128}(x-4)^2 + \frac{21}{259}(x-4)^3 + \frac{105}{1024}(x-4)^4 + ......"